40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jan 15 12:04:26 2010, Sanjit, mevans, Update, Adaptive Filtering, Canceling noise again! OAF_15JAN2010.png
    Reply  Tue Jan 26 19:51:44 2010, Sanjit, rana, Update, Adaptive Filtering, OAF details C1ASS_TOP.pngC1SUS_SRM_XYCOM1.pngUntitled.png
       Reply  Mon Feb 1 18:31:00 2010, Sanjit, Update, Adaptive Filtering, OAF details 
          Reply  Mon Feb 1 21:51:12 2010, Sanjit, Update, Adaptive Filtering, OAF details 
       Reply  Fri Feb 5 00:52:55 2010, Sanjit, Update, Adaptive Filtering, OAF at 0.1-1.0 Hz OAF_04FEB2010_noOAF.pngOAF_04FEB2010.png
          Reply  Fri Feb 5 01:04:58 2010, Sanjit, Update, Adaptive Filtering, OAF at > 5Hz OAF_04FEB2010_highFreq.png
             Reply  Sat Feb 6 00:10:08 2010, Sanjit, Update, Adaptive Filtering, OAF at > 5Hz 
Message ID: 2548     Entry time: Tue Jan 26 19:51:44 2010     In reply to: 2516     Reply to this: 2555   2571
Author: Sanjit, rana 
Type: Update 
Category: Adaptive Filtering 
Subject: OAF details 

We turned on the OAF again to make sure it works. We got it to work well with the Ranger as well as the Guralp channels. The previous problem with the ACC is that Sanjit and Matt were using the "X" channels which are aligned the "Y" arm. Another casualty of our ridiculous and nonsensical coordinate system. Long live the Right Hand Rule!!

The changes that were made are:

  • use of RANGER channel (with ACC_MC1_X and/or ACC_MC2_X)
  • mu = 0.01, tau = 1.0e-6, ntaps = 2000, nDown = 16
  • nDelay = 5 and nDelay = 7 both work (may not be so sensitive on delay at low frequencies)
  • Main changes: filter bank on the PEM channels - ASS_TOP_PEM_## filters: 0.1:0, 1:, Notch24, AA32, gain 1
  • Added the AI800 filter for upsampling in MC1 (should not matter)

Other parameters which were kept at usual setting:

  • CORR: AI32, gain = 1
  • EMPH: 0.001:0, AA32, gain = 1
  • ERR_MCL: no filters, gain = 1
  • SUS_MC1: no filter, gain = 1
  • PEM Matrix: All zero except: (24,1), (15,2), (18,3)
  • ADAPT path filter: union of CORR and EMPH filters, gain 1
  • XYCOM switches # 16-19 (last four on the right) OFF 

Screenshots are attached.

Burt snapshot is kept as: /cvs/cds/caltech/scripts/OAF/snaps/ass_burt_100126_211330.snap

taken using the script: /cvs/cds/caltech/scripts/OAF/saveOAF

we should put this in ASS screen.


ERROR Detected in filter ASS_TOP_PEM_24 (RANGER): 1: was actually typed as a 1Hz high pass filter!

(Correcting this one seems to spoil the adaptation)

Possibly this makes sense, we may not want to block witness signals in the 0.1-20 Hz range.


  11:40 PM: Leaving the lab with the OAF running on 5 PEM channels (Ranger + Guralp 1&2  Y & Z). There's a terminal open on op440m which will disable the OAF in ~2.8 hours. Feel free to disable sooner if you need the MC/IFO.

Attachment 1: C1ASS_TOP.png  33 kB  Uploaded Tue Jan 26 23:09:05 2010  | Hide | Hide all
C1ASS_TOP.png
Attachment 2: C1SUS_SRM_XYCOM1.png  33 kB  Uploaded Tue Jan 26 23:19:27 2010  | Hide | Hide all
C1SUS_SRM_XYCOM1.png
Attachment 3: Untitled.png  78 kB  Uploaded Tue Jan 26 23:37:40 2010  | Hide | Hide all
Untitled.png
ELOG V3.1.3-